OpenStudy (anonymous):

Expand (2^x + 3/y)^4 @MIT 18.02 Multiva…

OpenStudy (agreene):

\[\frac{81}{y^4}+\frac{27\times2^{x+2}}{y^3}+\frac{27\times 2^{2x+1}}{y^2}+\frac{3\times 2^{3x+2}{y}+2^{4x}\]

OpenStudy (anonymous):

umm.. i'm not sure if i'm reading this right. Is it...

OpenStudy (anonymous):

81y^4 + 27(2^x+2)*y^3 + 27(2^2x+1)*y^2 + 3(2^(3x+2))*Y

OpenStudy (agreene):

\[\frac{81}{y^4}+\frac{27\times2^{x+2}}{y^3}+\frac{27\times 2^{2x+1}}{y^2}+\frac{3\times 2^{3x+2}}{y}+ 2^{4x}\]

OpenStudy (anonymous):

oh, ok. Thx! :)